A box having a mass of 20 kg is suspended from two ropes, rop A and rope B, which are at the angles 37 degrees and 60 degrees respectively, with the horizontal. Findthe tension in each rope.
Answer #1
At equilibrium position, the net force acting on the system is zero.
T1sin37+T2sin60 = mg -------------- (1)
T1cas37 = T2cas60
T1 = 0.6260 T2 -------------- (2)
From equation (1)&(2),
0.3767 *T2 +0.8860*T2 = 20*9.8
T2 = 155.21 N
and
T1 = 97.16 N

answered by: brit